A line passes through the points (-10,-4) and (-1,2)
What is the y-intercept of the line?

Answer by stanbon(75887)

A line passes through the points (-10,-4) and (-1,2)
What is the y-intercept of the line?
-------------------------
Form:: y = mx + b
Solve for "m" and for "b".
-------
m = (2--4)/(-1--10) = 6/9 = 2/3
-------
Solve for "b":
y = (2/3)x + b
2 = (2/3)-1 + b
2 = -2/3 = b
b = 8/3
----
That is the y-intercept.
==========================
Cheers,
Stan H.
================